### v2

# Test-Time Efficient Pretrained Model Portfolios for Time Series Forecasting

Abstract
Is bigger always better for time series foundation models? With the question in mind, we explore an alternative to training a single, large monolithic model: building a portfolio of smaller, pretrained forecasting models. By applying ensembling or model selection over these portfolios, we achieve competitive performance on large-scale benchmarks using much fewer parameters. We explore strategies for designing such portfolios and find that collections of specialist models consistently outperform portfolios of independently trained generalists. Remarkably, we demonstrate that post-training a base model is a compute-effective approach for creating sufficiently diverse specialists, and provide evidences that ensembling and model selection are more compute-efficient than test-time fine-tuning.
1 Introduction
The dominant paradigm in pretrained time series models is the “bigger is better” view, supported by evidence that larger models improve forecast accuracy. Building on this observation, recent work has focused on scaling up both model size and training data to build better zero-shot forecasters (Das et al., 2024; Woo et al., 2024; Ansari et al., 2024a; Shi et al., 2024a; Edwards et al., 2025). However, these large, monolithic models come with high training and inference costs, which limits their practical use. To offset the need for ever-larger models, a promising alternative has emerged: allocating extra compute at test time. In NLP and computer vision, this includes generating multiple outputs for a single input and aggregating them—either by sampling from a single model (Sun et al., 2024) or using multiple models (Mavromatis et al., 2024). These examples, which partially inspire our work, belong to a broader class of methods that continue to achieve increasingly successful results across domains. Nevertheless, such strategies have not yet found their way to time series forecasting.
Inspired by these developments, we explore an alternative path. Instead of scaling up a single model, we propose to build a portfolio of smaller pretrained time series models and combine them at test time. This strategy offers greater flexibility and efficiency, but introduces two key challenges. First, how can we train a diverse set of models without incurring the full cost of training each one from scratch? Second, how should we combine the models at inference time to make accurate and efficient predictions? We address both challenges and show that this approach can match or even outperform larger monolithic models, while reducing inference costs and enabling modular design. More specifically, we make the following contributions.
  1. 1.
    We show that a portfolio of small pretrained models (based on encoder-decoder Chronos-Bolt Ansari et al. (2024b) architecture), each specializing on a subset of the training corpus, can match the accuracy of large monolithic models, while being more compute-efficient. By combining models at test time using ensembling or model selection, our approach achieves accuracy comparable to state-of-the-art pretrained forecasters, while dramatically reducing inference cost. We further observe that portfolios follow similar compute–performance scaling trends as monolithic models.
  2. 2.
    We introduce an efficient strategy for constructing diverse model portfolios without training each model from scratch. The approach starts by pretraining a generalist model on the full data distribution, followed by targeted fine-tuning on data partitions defined by metadata such as frequency and domain, obtaining specialists (see, Figure 1). This reduces overall training time by 10x and yields a portfolio that maintains diversity and delivers accurate forecasts.
  3. 3.
    We explore different methods of combining pretrained portfolios at test time. Unsurprisingly, we find that the performance of our portfolios depends on the ability to quickly identify the best sparse combination among them in a new test task. We find that performing model selection or greedy ensemble selection among specialists is an efficient approach when compared to other methods of utilizing compute at test time (e.g., with fine tuning).
Figure 1: A diagram depicting the approaches explored in this work. (a) Typical setup for pretraining, where a single large model is trained on a large corpus composed of many different datasets. At test time, the model can be invoked zero-shot as well as with finetuning at test-time and used to obtain forecasts. (b) Building a model portfolio by training an array of smaller models, where each specialist is trained on a single smaller corpus representing a different modality or domain. At test time, these models are combined via fitting an ensemble or by selecting the best model, and the selected combination (or, single model) is used for forecasting. (c) Specialists can also be trained in two-stages, by post-training the generalist model. We demonstrate that this approach for inducing diversity into the model portfolio results in comparable accuracy, while reducing the training-time compute by an order of magnitude. The approach also leads to the same forecast accuracy as (a), in return for a much smaller number of total parameters used actively for inference.
2 Background and Related Work
2.1 Time series forecasting
The objective of time series forecasting is to predict the future $H$ values of a time series $x$, given the previous $C$ observations. Formally, we consider a time series $x=[x_{1},\dots,x_{C}]$ where $x_{t}\in\mathbb{R}$ denotes the value of the time series $x$ at time step $t$. The goal of probabilistic forecasting is to approximate the distribution of future time series values $x_{C+1:C+H}$ given the history $x_{1:C}$ with a model $p_{\theta}$
$p_{\theta}(x_{C+1:C+H}|x_{1:C})\approx p(x_{C+1:C+H}|x_{1:C}).$
Forecasting approaches fall into three broad categories (Benidis et al., 2022; Januschowski et al., 2020). Local models, such as ETS and ARIMA (Hyndman and Athanasopoulos, 2018), train separate models, often fully determined by several parameters ($\sim$10), for each individual time series. Global models, including deep learning methods (Salinas et al., 2020; Rangapuram et al., 2018; Oreshkin et al., 2020; Nie et al., 2023), learn shared parameters ($\sim$100K) across a single dataset of multiple related time series (e.g., energy consumption of different households). More recently, pretrained models such as Chronos (Ansari et al., 2024a), TimesFM (Das et al., 2024), and Moirai (Woo et al., 2024) have adapted ideas from large language models to time series forecasting. These models are trained on large and diverse collections of time series data and aim to generalize to new forecasting tasks without task-specific fine-tuning. With parameter counts typically exceeding 10M, they serve as universal forecasters capable of making accurate zero-shot predictions across a wide range of domains. To reduce inference costs associated with such large models, Ekambaram et al. (2024) proposed Tiny Time Mixers (TTMs)—lightweight models with around 1M parameters, specialized for specific context and forecast settings. Our approach shares the same goal of reducing the test-time cost, but instead of specializing a single small model per setting, we train a portfolio of small models and select or combine them at test time.
2.2 Model and forecast combination
Traditional model combination methods, such as bagging (Breiman, 1996a) and boosting (Freund and Schapire, 1997), improve prediction accuracy by aggregating the outputs of multiple complementary models. Similar ideas have been applied in time series forecasting under the name forecast combination (Wang et al., 2023b). Forecasts are often combined using weighted averages of individual model predictions. These weights are typically shared across all time series in the dataset and are either set uniformly or in proportion to each model’s validation performance (Pawlikowski and Chorowska, 2020). More flexible strategies learn these weights directly from data, for example using greedy ensemble selection (Caruana et al., 2004; Shchur et al., 2023).
In the context of pretrained models, some works have explored combining multiple LLMs, such as Mavromatis et al. (2024), which studies independently trained base models and their ensembles at test time. Here, the authors compare methods including weighted ensembling, uniform ensembling, and model selection. While their work is closely related to ours in the context of foundation model portfolios, it focuses exclusively on language tasks. In contrast, we focus on probabilistic time series forecasting, where the use of model combination and more generally test-time computation remains largely unexplored. Furthermore, beyond comparing model combination strategies as in Mavromatis et al. (2024), we also investigate how to build a diverse portfolio of test-time efficient models.
Another related approach for pretrained models is the mixture-of-experts (MoE) framework (Jacobs et al., 1991; Shazeer et al., 2017), where component models and combination rules are jointly trained. This approach is shown to be effective in time-series forecasting tasks as well (Shi et al., 2024b; Liu et al., 2024). Although these works showed that MoE can help reducing inference time and cost, it limits flexibility. In contrast, our approach trains the component models in the portfolio independently, and allows the use of different fusion strategies at test-time. Notably, this flexibility enables the fusion rule to be adapted according to available computational resources during inference. Additionally, our approach enhances interpretability during the prediction by clarifying the contribution of each portfolio member, whose training procedures are transparent.
3 Chroma: A Portfolio of Small Pretrained Forecasters
In this section, we introduce our method for building portfolios of small pretrained forecasting models. Our approach consists of two stages. In the training stage, our goal is to build a diverse collection of small forecasting models. At test time, we combine their predictions to maximize accuracy for the given forecasting task. Applying this methodology to the publicly available pretrained models Chronos-Bolt (Ansari et al., 2024b), we construct our model portfolio, Chroma, which we describe in detail below.
3.1 Model architecture
We focus on portfolios composed of models with the same architecture, each pretrained independently on a different subset of the target distribution. The key difference is that we pretrain the full portfolio once, ahead of time, and reuse it across tasks—just as one would with a single large pretrained model.
We build Chroma using the training setup and architectures of Chronos-Bolt (Ansari et al., 2024b), a publicly available pretrained time series model based on the T5 encoder-decoder architecture (Raffel et al., 2020). Both pretraining and fine-tuning use average quantile loss, consistent with the original Chronos-Bolt implementation. Chronos-Bolt generates multi-step-ahead quantile forecasts and is trained on a diverse collection of time series datasets. We select this model due to its strong benchmark performance, open-source training pipeline and data, and support for a range of model sizes (e.g., tiny, mini), which enables flexible experimentation under different compute budgets. We train multiple versions of the model on different training datasets to induce diversity. We focus on smaller models ranging from 1M to 9M parameters, corresponding roughly to the tiny (9m) and below scale. Further implementation and training details are provided in Appendix B.
3.2 Building diverse portfolios
We now turn to the question of how to construct model portfolios that can be effectively combined at test time. It is well established that ensembles benefit from diversity among their members, provided each model is reasonably accurate (Zhou, 2012). However, diversity and individual model accuracy often trade off against each other (Wood et al., 2023).
In the context of pretrained models, diversity must be introduced during training, as weights are typically frozen at test time. The standard Chronos-Bolt training strategy aggregates all available time series datasets into a single, large corpus and trains one generalist model to perform well on average. In contrast, we aim to exploit structure in the data by explicitly training specialist (see, (Hinton et al., 2015)) models on disjoint subsets of the training corpus. Note that these models are not trained at once as in mixtures-of-experts or on the same data distribution but with perturbed samples as in bagging. Instead, we train each specialist model independently on a distinct partition of the corpus representing a different subproblem within the broader target distribution.
Specifically, we partition the data along metadata dimensions that reflect distinct characteristics of the time series. We focus on two such dimensions: frequency and application domain. Each specialist model is trained on a smaller subset of data defined by one of these attributes. For example, a domain-based portfolio may include an “energy” specialist trained only on energy-related datasets, while a frequency-based portfolio may contain an “hourly” specialist trained solely on hourly-resolution data. In each portfolio, we also include a generalist model of identical size trained on the full corpus for comparison. The training corpus of Chroma includes Chronos training data sets, plus some new ones for underrepresented frequencies and domains. The partitions and statistics of the datasets used are specified in Appendix C.
3.3 Reducing training cost through post-training
While forming portfolios, training all specialists from scratch on large datasets can result in significant computational cost. To mitigate this, we introduce diversity into the portfolio through post-training: we first train a generalist model of a given size, then fine-tune it briefly on different subsets of the training data to produce specialist models. This avoids the overhead of training each specialist model from random initialization.
This strategy is particularly important in our setup, where some data partitions may contain far fewer examples than others—for example, datasets with yearly resolution typically have far fewer observations than those with minutely resolution. In such cases, training from scratch may be inefficient or impractical. By leveraging post-training on a shared base model, we reduce overall training time and compute requirements while still matching the accuracy of portfolios trained independently. The benefits of post-training become even more critical in domains with larger models. While the largest time series models today are around 1 billion parameters, NLP models can reach up to 1 trillion. In such regimes, training each specialist from scratch is often infeasible, making post-training a practical and scalable alternative.
3.4 Forecast combination at test-time
Given a new time series dataset that was not seen by the portfolio members during training, we can utilize the models through one of two strategies: model selection or ensembling. In the first approach, we simply select the single best-performing model based on its accuracy on a validation set. In the second approach, we combine the predictions from multiple models using weighted averaging $\hat{y}_{\mathrm{ens}}=\sum_{m=1}^{M}w_{m}\cdot\hat{y}_{m}$, where the weights are determined using the ensemble selection algorithm (Caruana et al., 2004). For completeness, we provide the full description of the ensemble selection algorithm in Appendix D. Both strategies depend on validation sets produced through time series cross-validation (Hyndman and Athanasopoulos, 2018) to rank or weight the predictions in the model portfolio.
4 Experiments and Results
Setup. We extend the range of Chronos-Bolt model sizes to smaller variants and train portfolios of model sizes 1m, 2m, 4m, and tiny (9m). Details of these model architectures are provided in Appendix B. We extend the training corpus of Chronos [^1] https://huggingface.co/datasets/autogluon/chronos_datasets and partition this new corpus across two dimensions: frequency and domain.
For each model size, we first train a generalist using the original Chronos-Bolt corpus and a 200K-gradient-step training procedure following the reference implementations Ansari et al. (2024b), and incorporating mixup and synthetic data as specified in Ansari et al. (2024a). We then construct a portfolio of specialists by further fine-tuning (i.e., post-training) the generalist weights on different subsets of the training corpus for 1K gradient steps. In exploratory experiments, we also tested 3K, 10K, and 20K post-training steps, but observed no significant performance differences. As a result, we did not tune this variable further or include it as an experimental factor, and we use 1K gradient steps in our main experiments. This post-training phase amounts to only 0.5% of the generalist’s training time, making our method of constructing specialists extremely lightweight. Each portfolio includes the generalist and its associated specialists.
To evaluate a portfolio on a target task, we hold out the last $H$ observations from the training set to create a validation window. All models in the portfolio generate forecasts for this window. We then either select the model with the lowest loss on the validation set, or fit a weighted ensemble using the greedy ensemble selection algorithm (Caruana et al., 2004) implemented in AutoGluon (Shchur et al., 2023). These are based on rolling-window evaluations (also referred to as backtesting or out-of-time folds) and can be performed even with a single time series. It can also be performed by increasing the number of evaluation windows. Namely, instead of using a single window for selecting which specialists provide the best forecasts, one could rely on multiple. The selected model or ensemble is then applied to the test set to produce final forecasts. Note that we use the same methodology throughout all experiments.
Most of our results are reported on Chronos Benchmark II [^2] https://huggingface.co/spaces/autogluon/fev-leaderboard (BM2). Importantly, all datasets in the benchmark are zero-shot examples for our models (these datasets, including their training partitions, were not used in training of Chroma portfolios). Following the benchmark, we report the probabilistic forecasting results using the weighted quantile loss (WQL) and point forecasting results using the mean absolute scaled error (MASE). Finally, we also evaluate the larger Chroma variants on GIFT-Eval [^3] https://huggingface.co/spaces/Salesforce/GIFT-Eval (Aksu et al., 2024). For both benchmarks, we report aggregate performance using the geometric mean of relative errors. Specifically, for each dataset, we divide the model’s error by that of a baseline (Seasonal Naive unless stated otherwise), and then compute the geometric mean across datasets (Fleming and Wallace, 1986). Further details on the experiment setup are given in Appendix B.
4.1 Performance and test-time efficiency
Figure 2: Results on Chronos Benchmark II (BM2) and GIFT-Eval. Results reported are for probabilistic forecasting, with weighted quantile losses (WQL), scaled relative to Seasonal Naive model, aggregated across all data sets using geometric mean. For Chroma, Best refers to performing model selection on a validation set while Ensemble refers to the ensemble selection algorithm.
Benchmark performance compared to the state of the art.
We report overall results for Chroma on BM2 in Figure 2 (left), comparing the performance of 4m and tiny (9m) Chroma portfolios of frequency specialists to the state of the art in pretrained time series models. We find that, despite having as little as 4M active parameters at test time, the Chroma portfolio under best model selection performs comparably to much larger monoliths such as Moirai-1.1 Large (311M parameters) (Woo et al., 2024), TimesFM-2.0 (500M parameters) (Das et al., 2024) or Chronos-Bolt Base (205M parameters).
We also report results on the GIFT-Eval benchmark, in Figure 2 (right), including only pretrained architectures and statistical baselines for comparison. In rolling window evaluation tasks, we perform model selection or ensembling on the Chroma model portfolio on the first evaluation window alone, and infer with this fixed selection across the subsequent windows. That is, the ensemble was fixed in the training window of the first rolling window, and was not refit in subsequent windows. In this benchmark, too, Chroma’s aggregated probabilistic forecasting performance over 97 tasks is close to some much larger pretrained architectures, including TabPFN-TS (Hoo et al., 2025) which leverages significant test-time computation for in-context learning with a tabular foundation model. Note that we report results for TTM (Ekambaram et al., 2024) in zero-shot mode only. Moreover, in order not to misrepresent the performance of TTM, which specializes in high-frequency time series, we omit TTM’s results on BM2 which primarily contains low-frequency time series datasets. Further details on benchmark evaluations, including point forecasting performance, are given in Appendix G.
Portfolio design and combination method.
We now investigate the impact of key design choices in portfolio construction and test-time combination across different model sizes. To better isolate the effect of these choices, we report results on BM2 using a different reference point: instead of comparing to the Seasonal Naive baseline, we compute relative errors with respect to a single generalist model of size 1m. We also include a similarly sized portfolio of generalists (models trained with the same large corpus of data but with different random seeds).
We observe that building time series model portfolios by partitioning datasets by frequency is preferable to partitioning them by application domain, as this results in $\sim 5\%$ decrease in overall MASE and $\sim 4\%$ in WQL, across all model sizes considered (Table 1). This is consistent with earlier work such as TTM (Ekambaram et al., 2024) who used frequency to partition datasets. When comparing the use of ensemble vs. model selection, although ensemble selection leads to a slight improvement across model sizes and metrics, we find the evidence for this is not as strong as our previous conclusion. Therefore, while using ensembling to combine Chroma models can result in improvements in accuracy, this difference may not be strong enough to justify the added computational costs of using multiple models for inference.
Finally, we observe that simply combining multiple generalists trained using the same large data set does not result in significant performance gains over just using a single generalist. We hypothesize this is due to variance being a negligible factor in the test error committed by models pretrained on large corpora—an observation that can be supported by Lin et al. (2024), who argues that current LLMs remain underfitted owing to the relatively small number of training epochs. We discuss this observation further in Appendix F. Specifically, we show that in the regime where Chroma models are trained, bias component of the generalization error dominates the error due to variance. Since specialists reduce bias within a specific part of the overall data distribution, intelligent model combination or ensembling at test-time can reduce the overall bias of the portfolio. This is as opposed to the traditional ensembling approaches like bagging that tries to reduce error by decreasing variance.
Table 1: Aggregated relative WQL and MASE on BM2, scaled against a single generalist model of size 1m. The best two models for each metric and model size are given in bold.
WQL MASE
1m 2m 4m tiny 1m 2m 4m tiny
Domain Ensemble (Greedy) 0.957 0.936 0.932 0.915 0.964 0.932 0.928 0.918
Model Selection 0.963 0.950 0.939 0.923 0.973 0.948 0.933 0.930
Frequency Ensemble (Greedy) 0.926 0.898 0.890 0.896 0.910 0.884 0.878 0.887
Model Selection 0.918 0.916 0.880 0.909 0.920 0.899 0.887 0.893
Generalists Ensemble (Greedy) 0.987 0.951 0.939 0.919 0.974 0.944 0.931 0.928
Model Selection 0.990 0.966 0.942 0.944 0.979 0.961 0.935 0.941
Single Generalist None 1.000 0.977 0.960 0.958 1.000 0.971 0.962 0.961
Figure 3: Scaling behavior of Chroma portfolios. Results are presented for performing model selection from a Chroma portfolio of domain or frequency specialists. Each individual run is for an independently trained generalist in the portfolio, with four trials reported per experiment setting. $\alpha$ refers to the slope of the scaling fit, and asterisks denote statistical significance of the fitted coefficient at the 5% level (p < 0.05). Reported results are aggregated across BM2.
Scaling behavior.
Having observed that our conclusions hold across different model sizes, we turn our attention to quantifying how Chroma scales. Similar to recent works (Hoffmann et al., 2022; Kaplan et al., 2020; Shi et al., 2024a; Edwards et al., 2025), we carry out ordinary least squares fits to the test error of Chroma (on BM2) and the number of active parameters at inference time. We present results for model selection on the Chroma portfolio in Figure 3. We observe that despite individually training on much smaller datasets, the overall model portfolio’s performance scales at rates comparable to the individual generalist models. In further experiments, we found that our training setup of post-training on generalists to obtain the portfolio is essential to this conclusion (see, Appendix G). We therefore conclude that, provided datasets can be proportionally scaled, our approach can be scaled up to larger model sizes to obtain even better performance.
Test-time compute efficiency.
Chroma delivers strong forecast accuracy, but at the cost of additional test-time computation. Since we must either select the best model from the portfolio or combine multiple models via greedy ensembling, each model in the portfolio must run inference on both the validation and test sets. This raises a natural question: how does this compute overhead compare to alternative strategies for adapting pretrained models at test time?
The simplest option is to use a single generalist model in a zero-shot manner—no adaptation, and only one forward pass. Our approach sits in the middle: it avoids fine-tuning but requires multiple inference passes and some lightweight model selection or combination logic. At the high end of the compute spectrum is test-time fine-tuning, where a single generalist model is adapted to the target task using gradient updates.
To quantify these trade-offs, we compare all three strategies—zero-shot generalist, Chroma (with model selection or ensembling), and fine-tuned generalist—by measuring total test-time floating point operations (FLOPs) and reporting accuracy across the board. Fine-tuning is done with 1K gradient steps for reference. This lets us position Chroma in terms of both accuracy and compute efficiency relative to common adaptation strategies. Our results are presented in Figure 4, where different model sizes and combination methods are shown along the efficient frontier of accuracy vs. the total test-time compute required. Fine-tuning a generalist model for 1K gradient steps yields strong accuracy but requires nearly 10x more compute than simple forward passes. Chroma offers a middle ground: it achieves competitive accuracy using a fixed portfolio of small models, with only lightweight model selection or ensembling at test time. This places it near the center of the efficiency frontier, offering a favorable trade-off between accuracy and computational cost. Further fine-tuning may improve accuracy, but it shifts the operating point toward significantly higher compute requirements.
[…]
4.2 Further experiments
Ablation studies.
We perform three ablation studies summarized in Table 2. A key element in the portfolios we consider is the inclusion of a generalist (see, e.g., Hinton et al. (2015)). We estimate the effect of excluding the generalist on MASE and WQL errors averaged across different model sizes, specialist types (domain vs. freq) and datasets. In both cases, we find average error increases of $\sim 2\%$, while these regressions in performance are not measured to be statistically significant.
Table 2: Summary results of ablation studies. Results are given in the estimated increase in normalized error (where seasonal naive error is 1) on BM2 for the ablation performed, and the p-value associated with the t-test of the regression coefficient.
MASE WQL
Increase in Error p-value Increase in Error p-value
No Generalists in Portfolio 0.016 0.107 0.017 0.118
Ensemble - Performance-weighted 0.186 0.000 0.142 0.000
Ensemble - Simple Average 0.242 0.000 0.180 0.000
Specialists Trained From Scratch 0.002 0.798 -0.001 0.918
Second, we observe that the benefit of specialist portfolios comes primarily from the diversity they introduce. However, this diversity is only useful if we combine models intelligently at test time. In particular, our ensembles do not work simply by reducing variance through averaging, as might be expected in settings where randomness in training dominates model differences. We confirm this in ablation studies: both simple averaging and performance-weighted averaging (cf. Pawlikowski and Chorowska, 2020) lead to significant performance drops when applied to our pretrained portfolios. This contrasts with much of the prior forecasting literature, where ensembles of task-specific models are often effective even with naive combination strategies (Wang et al., 2023b).
An essential part of our methodology in training a diverse portfolio of models is to pretrain a single generalist before inducing diversity by post-training with different data subsets. In our final ablation study, we compare this to training all specialists from random initializations (see, Figure 1 (b)). On average, this does not lead to any regressions in performance. However, in scaling studies we find that post-trained models scale better (see, Appendix G).
Credit assignment and interpretability of the portfolio.
We also investigate which specialists are activated during ensembling or model selection for a given test task. Figure 5 illustrates this activation pattern. The heatmaps are structured as matrix plots, where rows represent tasks and columns represent specialists (categorized by domain or frequency). We observe that all specialists are activated for some tasks, indicating no “mode collapse” where only a single model dominates. Notably, when the test task shares a domain or frequency with a particular specialist, that specialist tends to be highly activated. This suggests a practical efficiency gain: in many real-world applications, the nature of the test-time task is known in advance. This knowledge can be leveraged to manually select relevant specialists, thereby reducing test-time computational cost. Additionally, Figure 5 also offers a layer of interpretability. For each task, the specialist weights provide meaningful insights that can aid practitioners in understanding the model’s behavior.
Figure 5: Distribution of ensemble weights for 4m specialist portfolios across distinct tasks, grouped with respect to the their domain or frequency information.
5 Conclusion
In this work, we introduced Chroma: a portfolio of small, pretrained time series forecasters. Chroma consists of specialist models, each fine-tuned on disjoint subsets of the training data from a generalist pretrained model. When model selection or ensembling is applied at test time, Chroma achieves competitive performance on forecasting benchmarks from Ansari et al. (2024a) and Aksu et al. (2024), matching the accuracy of much larger monolithic models. Our experiments demonstrate the test-time efficiency of Chroma, as well as its scalability. We believe that the methodology behind Chroma opens a promising new direction for enhancing the accuracy of pretrained forecasting models, especially under limited compute budgets. Furthermore, while our focus is on time series forecasting, the underlying principles may also extend to other domains such as natural language processing and computer vision. Our strategy offers an alternative to approaches like Best-of-$N$ which samples from a single base model. Instead, it proposes an inference-aware portfolio formation strategy: train a single generalist base model, fine-tune it into a portfolio of specialized experts, and apply model selection at test time. Moreover, we focused on test-time efficiency because inference often dominates the computational budget in deployed forecasting systems.
Limitations and future work.
Although we did not study the impact of the pretraining phase in detail, our design choices follow standard practice: T5-architecture based Chronos-Bolt model with straightforward modifications, see Appendix B. Also, our theoretical observations in Appendix F provides a bias-variance perspective that helps explain the observed gains and suggests implications for the role of pretraining. Understanding how the number of training iterations and other pretraining hyperparameters interact with our conclusions would be a valuable direction for future research. Similarly, our experiments were conducted with encoder-decoder transformer architectures. Future work can investigate portfolios of transformers based on other types of architectures. Moreover, while we hand-crafted the specialist portfolios using known dataset features (e.g., domain or frequency), future work could explore more principled approaches. For instance, foundation model specialists can be formed by incorporating boosting-style objectives (Freund and Schapire, 1997). Finally, this work focused on weighted ensembling; future research could explore more sophisticated methods such as stacked ensembling (Breiman, 1996b) that can lead to higher accuracy.
APPENDIX
Appendix A Additional Literature
Prior work shows that ensembles of neural networks outperform single networks and provide better uncertainty quantification. Specifically, it has been reported that training the same model architecture from different random initializations is an effective strategy for sufficiently diverse portfolios Lee et al. (2015), and deep ensembles utilize this idea Lakshminarayanan et al. (2017); Fort et al. (2019); Zaidi et al. (2021). At the foundation-model scale, however, such work is less common. Notably, Sun et al. (2022) builds LLM ensembles by varying random seeds, while Wang et al. (2023a) uses low-rank adapters for resource efficiency. We adopt a similar approach while forming our generalist portfolios and likewise observe gains over single models. However, we further show that larger improvements can be achieved with specialist portfolios, where models are fine-tuned on targeted subsets of the data distribution. As discussed in Appendix F, in the foundation-model regime, variance reduction from ensembling is less decisive than improved coverage of heterogeneous data.
In time-series context, we also find that partitioning by frequency to induce specialization across portfolio members performs particularly well. This aligns with a broader line of work that exploits frequency-based representation and basis expansions for time-series forecasting. Namely, Darlow et al. (2024) proposes a foundation model that predicts coefficients of sinusoidal bases for functional decomposition, and Kim et al. (2024) represents time series directly in the frequency space. As opposed to those methods aiming to learn a single, monolithic model, our methodology is focused on specialization, i.e., training multiple experts on frequency-based clusters and aggregating their predictions. In classical statistical forecasting, beyond our baselines such as Auto-ETS and Auto-ARIMA, there are also recent methods such as FFORMA (Montero-Manso et al., 2020), which learns to predict ensemble weights over a pool of univariate forecasters using time-series features, and SCUM (Petropoulos and Svetunkov, 2020), which aggregates established statistical methods via median-based fusion.
Appendix B Details on Model Architectures and Hyperparameters
We adopt Chronos-Bolt as the architecture for Chroma (Ansari et al., 2024b). Chronos-Bolt is based on the T5 encoder-decoder architecture (Raffel et al., 2020). It chunks the historical time series context into patches of multiple observations, as in PatchTST (Nie et al., 2023), which are then input into the encoder. The decoder then uses the representations generated by the encoder to directly generate quantile forecasts across multiple future steps—a method known as direct multi-step forecasting. This differs from the original Chronos (Ansari et al., 2024a) models that rely on autoregressive decoding. The Chronos-Bolt architecture following a standard pattern — namely, patching the inputs, using transformer blocks, and multi-step-ahead quantile prediction — makes it a representative choice for time series foundation models, as this pattern is similar to other recent SOTA time-series models.
We also adopt the training methodology of Chronos-Bolt, which was trained on nearly 100B time series observations from varying frequencies and application domains. The training objective is the mean weighted quantile loss (WQL) over 9 equally spaced quantiles between $0.1$ and $0.9$. For simplicity, we keep the same training corpus as Chronos-Bolt when training our generalist models. When training specialists, we added new datasets to the training corpus to balance data for underrepresented frequency and domain classes (see, Appendix C. Our generalist models are not trained on these new datasets to remain consistent with the Chronos-Bolt setup. Although we do not directly compare generalists and their corresponding specialists in our main experiments, we also explored training generalists on the full specialist dataset. We have observed that this had little effect. The performance of individual generalists, test-time fine-tuned generalists, and generalist portfolios considered in the main paper remained largely unchanged.
[…]
Table 3: Architectural hyperparameters for the model sizes considered in the paper. $d_{\text{ff}},d_{\text{kv}},$ and $d$ refer to the feedforward model hidden dimension, key-value dimension, and the model (embedding) dimension respectively. $L_{e}$ and $L_{d}$ denote the number of layers in the encoder and decoder.
$d_{\text{ff}}$ $_{\text{kv}} dd$ $L_{d}$ $L_{e}$
tiny 1024 64 256 4 4
4m 768 64 192 3 3
2m 640 32 160 2 2
1m 512 32 128 1 2
We trained and evaluated all models on NVIDIA A10G Tensor Core GPUs, used through AWS EC2 G5 family of instances. For reference, our 1m, 2m, 4m, and 9m (tiny) model sizes take $1.4$, $1.7$, $3.3$, and $4.8$ hours on average to train from scratch for 200K iterations, on a single A10G GPU. Post-training runs typically take several minutes.
Appendix C Details on Specialists
In order to train the models of Chroma portfolios, datasets inside the general training corpus [^5] https://huggingface.co/datasets/autogluon/chronos_datasets/ as well as additional datasets from Monash (Godahewa et al., 2021) and UCI (Asuncion et al., 2007) repositories, are partitioned in two ways, into six partitions of domain and five partitions of frequency. Details of these assignments, as well as some summary statistics of the data sets are provided in Table 4.
Models trained from scratch follow the exact same training procedure as the generalist models, including 200K training iterations, albeit with access to only a specific part of the training corpus.
For Chroma specialists, which were post-trained, we randomly selected a generalist model as the initial weights and fine-tuned each specialist via 1K gradient steps, using the same training setup as mentioned above. Due to computational constraints, we trained only one set of specialists for each model size and partitioning scheme, however during experiments, we evaluated each portfolio 5 times varying the generalist in the portfolio and reported their average. Similarly, for all other results reported on single generalists, we give an average of 5 independently trained models.
Table 4: Datasets used in training, summary statistics and and their assignments to domain and frequency partitions.
Nr. Time Series Nr. Observations Missing Rate Avg. Length Frequency Domain
indian_power_generation 13 13,858 0.0% 1066 daily energy
wind_farms_daily 337 119,549 16.8% 355 daily energy
iowa_liquor_subset 847 1,803,263 0.0% 2129 daily retail
online_retail_I 1,742 587,893 0.0% 337 daily retail
online_retail_II 4,772 2,369,710 0.0% 497 daily retail
monash/covid_mobility 450 180,379 13.0% 401 daily transport
monash/covid_mobility_without_missing_values 450 180,379 0.0% 401 daily transport
uber_tlc_daily 262 47,422 0.0% 181 daily transport
m4_daily 4,227 10,023,836 0.0% 2371 daily various
monash/bitcoin 18 82,458 9.8% 4581 daily various
monash/bitcoin_without_missing_values 18 75,364 0.0% 4187 daily various
monash/us_births 1 7,305 0.0% 7305 daily various
monash/sunspot 1 73,924 4.4% 73924 daily weather
monash/sunspot_without_missing_values 1 73,924 0.0% 73924 daily weather
monash/temperature_rain 422 305,950 0.5% 725 daily weather
ushcn_daily 1,218 47,080,115 7.7% 38654 daily weather
monash/electricity_hourly 321 8,443,584 0.0% 26304 hourly energy
solar_1h 5,166 45,254,160 0.0% 8760 hourly energy
wind_farms_hourly 337 2,869,414 17.0% 8515 hourly energy
mexico_city_bikes 494 38,687,004 0.0% 78314 hourly transport
monash/pedestrian_counts 66 3,132,346 0.0% 47460 hourly transport
monash/rideshare 156 84,396 44.2% 541 hourly transport
taxi_1h 2,428 1,794,292 0.0% 739 hourly transport
uber_tlc_hourly 262 1,138,128 0.0% 4344 hourly transport
m4_hourly 414 373,372 0.0% 902 hourly various
monash/kdd_cup_2018 270 2,942,364 17.1% 10898 hourly weather
monash/kdd_cup_2018_without_missing_values 270 2,942,364 0.0% 10898 hourly weather
electricity_15min 370 41,936,458 0.0% 113342 subhour energy
monash/aus_solar_1min 1 493,149 0.0% 493149 subhour energy
monash/aus_wind_1min 1 493,144 0.0% 493144 subhour energy
monash/elecdemand 1 17,520 0.0% 17520 subhour energy
monash/london_smart_meters 5,560 166,528,896 0.0% 29951 subhour energy
monash/solar_10_minutes 137 7,200,720 0.0% 52560 subhour energy
monash/solar_4_seconds 1 7,397,222 0.0% 7397222 subhour energy
solar 5,166 543,049,920 0.0% 105120 subhour energy
taxi_30min 2,428 3,589,798 0.0% 1478 subhour transport
heart_rate 4 5,500 0.0% 1375 subhour various
monash/wind_4_seconds 1 7,397,147 0.0% 7397147 subhour weather
m4_monthly 48,000 11,246,411 0.0% 234 supra various
monash/m3_other 174 13,325 0.0% 77 supra various
brazil_gas_prices 187 128,471 0.0% 687 weekly energy
monash/electricity_weekly 321 50,076 0.0% 156 weekly energy
monash/solar_weekly 137 7,124 0.0% 52 weekly energy
m4_weekly 359 371,579 0.0% 1035 weekly various
monash/kaggle_web_traffic_weekly 145,063 16,537,182 0.0% 114 weekly web
weatherbench_daily 225,280 78,991,953,920 0.0% 350639 daily weather
wiki_daily_100k 100,000 274,100,000 0.0% 2741 daily web
Appendix D Details on Model Selection
Ensemble selection algorithm Caruana et al. (2004) is a standard approach for forecast combination in both point and probabilistic forecasting (Deng et al., 2022; Shchur et al., 2023). Given a set of models producing predictions $\hat{y}_{1},\ldots,\hat{y}_{M}$, where each $\hat{y}_{m}$ is the tensor of predictions for all item, time steps and quantile levels, the goal is to find optimal weights $w_{1},\ldots,w_{M}$ for the ensemble prediction $\hat{y}_{\mathrm{ens}}=\sum_{m=1}^{M}w_{m}\cdot\hat{y}_{m}$.
The ensemble selection algorithm optimizes validation loss by greedily adding models to an equally-weighted ensemble with replacement. Since an equally-weighted ensemble with replacement is equivalent to a weighted average with fractional weights, this can be interpreted as optimizing weights $w_{1},\dots,w_{M}$ via coordinate-wise ascent.
Algorithm: Greedy Ensemble Weighting
  1. 1.
    Initialize weights $w^{(0)}=\mathbf{0}\in\mathbb{R}^{M}$.
  2. 2.
    For iterations $j=1,\ldots,S$:
    1. (a)
      Select model that minimizes validation loss:
      $m^{(j)}=\mathop{\mathrm{arg\,min}}\limits_{m}\;L(\hat{y}_{\mathrm{ens}},y_{\mathrm{val}}),$
      where $\hat{y}_{\mathrm{ens}}$ uses the updated weights
      $\frac{(j-1)w^{(j-1)}+e_{m}}{j},$
      and $e_{m}$ is the $m$-th canonical basis vector.
    2. (b)
      Update weights:
      $w^{(j)}=\frac{(j-1)w^{(j-1)}+e_{m^{(j)}}}{j}.$
  3. 3.
    Return final weights $w^{(S)}$.
In our evaluation we use $S=100$ steps of ensemble selection and use the last window of $H$ observations in the training portion of each series as the validation set $y_{\mathrm{val}}$. As in the rest of the paper, we use the weighted quantile loss (WQL) as the loss function $L$.
Appendix E FLOPs Computation
In our test-time compute we use approximations to the total number of floating point operations used at test time. For this, we use approximations which we define here for completeness. As in Appendix B, let $L_{e},L_{d}$ denote the number of layers in the encoder and decoder, and $d$ the model dimension respectively. We also denote $T$ as the effective sequence length, which is the number of tokens that are given to the encoder as the context. For our work, $T=\frac{2048}{16}=128$ where 2048 is the original sequence length and 16 is the patch size, for all models considered. We compute
$\text{FLOPs}_{\text{forward}}~\approx~(L_{e}+L_{d})\times T\times d^{2}\times\Bigl(24+\dfrac{4T}{d}\Bigr).$
When model fine-tuning using backward passes are required at test time, we calculate
$\text{FLOPs}_{\text{train}}~\approx~3\times\text{FLOPs}_{\text{forward}},$
since the backward pass usually incurs about twice the cost of the forward pass, plus some overhead.
Note we compute two numbers for test-time compute, namely, total test-time compute and amortized compute. By the former, we refer to the total effort of performing model selection or ensembling with a given portfolio of $N$ models, and performing one inference pass with the resulting selection. For example, total test time compute for performing model selection is given as $(N+1)\times\text{FLOPs}_{\text{forward}}$ where the $N$ term results from the necessity to infer with each model once during selection, and the $1$ term to the single inference pass with the selected model. Similarly, for ensembling, this factor is $\approx(N+2.5)$ where $2.5$ is the average number of models selected for the ensemble. Finally, when we report amortized test-time compute, we omit the $N$ factor as the initial cost of performing the model selection or fine-tuning will be amortized over all future inference passes using the resulting ensemble or model.
Appendix F Bias–Variance in Foundation Model Portfolios
As we have seen in Table 1, portfolios composed of generalist models, each trained on the full training distribution, yield only small performance improvements over a single generalist model. In contrast, specialist portfolios in Chroma, whether based on frequency or domain, offer substantial gains. This may seem counterintuitive, since each generalist model is trained on the entire training distribution rather than a specific subset, and on average, a model from the generalist portfolio outperforms a model from the specialist portfolio in a direct, head-to-head comparison. In this section, we offer an explanation for this observation using the bias–variance decomposition of the expected generalization error.
The generalization error of a model is known to be decomposable as (Vapnik, 1999)
Average Test Error = Irreducible noise + Bias + Variance.
Here, irreducible noise is inherent to the task and cannot be reduced. Typically, model combination methods target either the bias or variance components to lower the overall error.
In a generalist portfolio, all models are trained on the same training corpus and therefore share same biases. As a result, ensembles’ overall bias remains unchanged from that of the individual models in the portfolio. However, such ensembles reduce variance. Indeed, if the models are independently sampled, the variance decreases by a factor of $N$, where $N$ is the number of models in the ensemble. Note that this setup is analogous to classical bagging in ensemble learning (Breiman, 1996a; Zhou, 2012), where multiple models trained on bootstrapped subsets of the full data set are averaged to reduce variance without affecting bias.
We argue that in the regime where pretrained time series models are trained, generalization error due to bias dominates the error due to variance. As a result, having a generalist portfolio does not result in significant performance gains over a single model.
To confirm this argument, we estimate the bias and variance components of the error of a single model. To do so, we train ten distinct generalists using different random seeds following the same training procedure outlined in Appendix B. In order to discard the “irreducible noise” component of the error, we test our models on synthetic signal inputs without a noise component. Specifically, we draw 10 thousand time series from a mixture of periodic Gaussian process kernels and offset these with random trends and constants. On this dataset, we compute
$\text{Bias}(x)=\left(\bar{f}(x)-y\right)^{2}\qquad\text{Variance}(x)=\frac{1}{M}\sum_{i=1}^{M}\left(f_{i}(x)-\bar{f}(x)\right)^{2},$
where $f_{i}(x)$ denotes the point forecast produced by model realization $i$ for test input $x$, $\bar{f}(x)$ is the average forecast across all $M$ realizations, and $y$ represents the true value of the signal corresponding to $x$ over the prediction horizon. The results are provided in Table 5. Observe that over all model size spectrum we consider in this work, the bias component of the error dominates the variance component across all generalist models.
Table 5: Bias and variance components of the error. It can be seen that bias component dominates the variance component across all model sizes.
Bias Variance
1m 65.305 10.067
2m 49.031 12.298
4m 22.009 6.423
tiny 20.063 8.367
Our observation for time-series foundation models aligns with the argument of Lin et al. (2024), which states that, in current large language models, the variance component of the error becomes strictly smaller than the bias because these models are typically trained for only one or a few epochs on their massive training corpora, leaving them underfit.
These observations also help us understand why Chroma works so well. While building pretrained model portfolios, Chroma trains specialist models by fine-tuning a generalist base model. Each of these specialists reduces bias within a specific part of the overall data distribution. As a result, a specialist portfolio consists of models with distinct biases.
Crucially, we leverage this diversity in the biases effectively through an intelligent model combination at test time. In contrast to traditional ensembling approaches, such as bagging, which reduce variance by averaging out randomness across models, Chroma’s advantage does not rely on variance reduction. Instead, it operates through targeted bias reduction. For instance, when applying model selection, it selects the most appropriate specialist for each test input, and the overall bias of the ensemble effectively becomes equal to the minimum bias among the models in the portfolio.
Table 2 supports this perspective, as well. A simple, uniform averaging over specialist portfolios fail to deliver performance gains comparable to model selection or greedy ensembling. This finding stands in contrast to much of the forecast combination literature, where heterogeneous ensembles of task-specific models typically benefit from averaging. In our case, however, the models are used in a zero-shot manner, and require test time input specific selection to be effective. This is consistent with our earlier observation that, in the pretrained regime, generalization error is dominated by bias rather than variance. Since variance is already low, reducing bias through targeted specialist selection offers a more effective way of improving performance.
Appendix G Additional Experiment Results
This section complements Section 4 by providing additional results and plots.
G.1 Scaling behavior
As discussed in Section 4, we perform ordinary least squares fits on the test error of BM2 to analyze the scaling behavior of Chroma. Each run corresponds to an independently trained generalist within the portfolio, with four trials reported per experimental setting. The parameter $\alpha$ denotes the slope of the scaling fit, and asterisks indicate statistical significance of the fitted coefficient at the 5% level (p < 0.05).
Figures 6 and 7 show the test error plotted against the number of active parameters at inference time for model selection and ensembling on the Chroma portfolio, respectively. Indeed, the overall portfolio performance scales at rates comparable to those of the individual generalist models.
In contrast, Figures 8 and 9 present results when the Chroma portfolio is trained from scratch, rather than post-training. Here, as mentioned in the main paper, the behavior differs significantly. Therefore, we conclude that post-training generalist models to form the portfolio is essential for Chroma’s scaling behavior.
We also provide scaling plots in terms of amortized test-time compute in Figures 10111213, calculated according to Appendix E, which provide the same conclusions.
Figure 6: Scaling behavior of Chroma portfolios. Results are presented for performing model selection from a Chroma portfolio of domain or frequency specialists, where each of these specialists are formed with post-training.
Figure 7: Scaling behavior of Chroma portfolios. Results are presented for performing ensembling on a Chroma portfolio of domain or frequency specialists, where each of these specialists are formed with post-training.
Figure 8: Scaling behavior of portfolios, when each of the specialists are trained from scratch. Results are presented for performing model selection from a portfolio of domain or frequency specialists.
Figure 9: Scaling behavior of portfolios, when each of the specialists are trained from scratch. Results are presented for performing ensembling on a portfolio of domain or frequency specialists.
Figure 10: Scaling behavior of Chroma portfolios with respect to the amortized test time compute, when the specialists are formed with post-training and model selection is performed during test-time.
Figure 11: Scaling behavior of Chroma portfolios with respect to the amortized test time compute, when the specialists are formed with post-training and greedy ensembling is performed during test-time.
Figure 12: Scaling behavior of portfolios with respect to the amortized test time compute, when the specialists are trained from scratch and model selection is performed during test-time.
Figure 13: Scaling behavior of portfolios with respect to the amortized test time compute, when the specialists are trained from scratch and greedy ensembling is performed during test-time.
G.2 Test-time compute efficiency
Figures 14 and 15 demonstrate the tradeoff between total test-time compute (in terms of FLOPs) and performance. FLOPs are estimated as described in Appendix E, and incorporate the costs of model fitting, fine-tuning, or model selection, in addition to a single forward pass for inference per item in the test set.
[…]
The conclusion remains the same in both figures. Chroma portfolios with frequency specialists lie at the test-time compute-efficiency frontier, whether the specialists are formed with post-training or trained from scratch.
Figure 14: Total test-time compute (GFLOPs) and performance tradeoff, presented for Chroma portfolios built through post-training.
Figure 15: Total test-time compute (GFLOPs) and performance tradeoff, presented for portfolios when the specialist models are trained from scratch.
G.3 Credit assignment and interpretability of the portfolio
In Figures 16, 17, 18, and 19, we extend the analysis from Figure 5 by applying the same methodology to portfolios composed of model architectures 1m, 2m, 4m, and tiny, as depicted in Appendix B. We examine which specialists are activated during ensembling for each test task. Like in Figure 5, the heatmaps are presented as matrix plots, with rows representing tasks and columns representing specialists (categorized by domain or frequency). Tasks are grouped according to their domain or frequency, and the weights are aggregated across tasks and normalized per group. We can see that our findings from Figure 5 remain consistent across portfolios of different model sizes.
Figure 16: Distribution of ensemble weights for 1m specialist portfolios across distinct tasks.
Figure 17: Distribution of ensemble weights for 2m specialist portfolios across distinct tasks.
Figure 18: Distribution of ensemble weights for 4m specialist portfolios across distinct tasks.
Figure 19: Distribution of ensemble weights for tiny specialist portfolios across distinct tasks.
G.4 Ablation Study
We perform three sets of ablations in this study, quantifying the effects of dropping generalists from ensembles, using simple averaging as the ensembling strategy, and for training Chroma portfolio models from scratch instead of post-training. For each study, we fit mixed linear models controlling for the evaluation dataset, and quantify the effect of the ablation and its statistical significance. While overall results for ablations are given in the main text, we also provide detailed results per portfolio type and model size as well.
Table 6: Ablation study, fixed effects and p-values for removing the generalist from model selection or ensembles for all model sizes and portfolio types.
Group WQL - Fixed Effect WQL - p-value MASE - Fixed Effect MASE - p-value
Portfolio Type domain 0.025 0.000 0.024 0.000
Portfolio Type freq 0.009 0.015 0.007 0.0536
Model Size 1m 0.017 0.015 0.012 0.184
Model Size 2m 0.017 0.015 0.018 0.0238
Model Size 4m 0.018 0.01225 0.018 0.014
Model Size tiny 0.016 0.0023 0.014 0.014
Overall N/A 0.017 0.000 0.016 0.000
Table 7: Ablation study, fixed effects and p-values for replacing greedy weighted averages with simple and performance weighted ensembles.
Group Ensemble Type WQL - Fixed Effect WQL - p-value MASE - Fixed Effect MASE - p-value
Portfolio Type domain Perf 0.149 0.000 0.193 0.000
Portfolio Type domain Simple 0.192 0.000 0.264 0.000
Portfolio Type freq Perf 0.136 0.000 0.179 0.000
Portfolio Type freq Simple 0.167 0.000 0.221 0.000
Model Size 1m Perf 0.142 0.000 0.179 0.000
Model Size 1m Simple 0.177 0.000 0.231 0.000
Model Size 2m Perf 0.137 0.000 0.179 0.000
Model Size 2m Simple 0.172 0.000 0.232 0.000
Model Size 4m Perf 0.152 0.000 0.199 0.000
Model Size 4m Simple 0.196 0.000 0.265 0.000
Model Size tiny Perf 0.138 0.000 0.186 0.000
Model Size tiny Simple 0.174 0.000 0.241 0.000
Overall N/A Perf 0.142 0.000 0.186 0.000
Overall N/A Simple 0.180 0.000 0.242 0.000
Table 8: Ablation study, fixed effects and p-values for replacing post-training with training each specialist from scratch.
Group WQL - Fixed Effect WQL - p-value MASE - Fixed Effect MASE - p-value
Portfolio Type domain -0.008 0.08575 -0.005 0.20533
Portfolio Type freq 0.007 0.08575 0.008 0.014
Model Size 1m -0.002 0.791 -0.003 0.554
Model Size 2m -0.009 0.08575 0.004 0.4732
Model Size 4m -0.001 0.791 0.009 0.1925
Model Size tiny 0.010 0.08575 -0.004 0.392
Overall N/A -0.001 0.791 0.002 0.554
G.4.1 Portfolios trained from scratch
In the result Table 2 of our ablation study, we show that there is no statistically significant difference in accuracy between portfolios trained from scratch and portfolios trained with post-training. In Table 9, we support this result by a rework of Table 1 that also includes portfolios of models that are trained from scratch, as opposed to post-training.
Table 9: Aggregated relative WQL and MASE on BM2 including portfolios trained from scratch, scaled against a single generalist model of size 1m.
WQL MASE
1m 2m 4m tiny 1m 2m 4m tiny
Domain Ensemble (Greedy) 0.957 0.936 0.932 0.915 0.964 0.932 0.928 0.918
Model Selection 0.963 0.950 0.939 0.923 0.973 0.948 0.933 0.930
Frequency Ensemble (Greedy) 0.926 0.898 0.890 0.896 0.910 0.884 0.878 0.887
Model Selection 0.918 0.916 0.880 0.909 0.920 0.899 0.887 0.893
Domain (scratch) Ensemble (Greedy) 0.953 0.927 0.923 0.911 0.947 0.924 0.925 0.905
Model Selection 0.987 0.928 0.943 0.919 0.968 0.936 0.942 0.919
Frequency (scratch) Ensemble (Greedy) 0.907 0.905 0.893 0.910 0.906 0.895 0.882 0.880
Model Selection 0.935 0.914 0.885 0.940 0.920 0.913 0.895 0.896
Generalists Ensemble (Greedy) 0.987 0.951 0.939 0.919 0.974 0.944 0.931 0.928
Model Selection 0.990 0.966 0.942 0.944 0.979 0.961 0.935 0.941
Single Generalist None 1.000 0.977 0.960 0.958 1.000 0.971 0.962 0.961
It can be seen that there is no significant difference between these two type of portfolios in terms of accuracy. Similarly, from the individual tasks point of view, we also were not able draw a direct conclusion about how from-scratch and post-trained models compare. Despite this performance similarity, the portfolios trained from scratch required 200x more compute for each specialist.
G.5 Detailed Benchmark Results
While we provide the aggregated results in the main text, in this section, we provide detailed results for each individual dataset included in the benchmarks. Specifically, for BM2, we report the results for WQL and MASE in Tables 10 and 11, respectively. For GIFT-Eval, WQL results are shown in Tables 12 and 13, while MASE results are presented in Tables 14 and 15.
Table 10: Weighted quantile losses (WQL) on Chronos Benchmark II
Chroma (4m, freq, best) Chroma (4m, freq, ens.) Chroma (tiny, freq, best) Chroma (tiny, freq, ens.) Chronos Bolt Base Chronos Bolt Mini Chronos Large Moirai Large TimesFM 2.0 Auto ARIMA Auto ETS Seasonal Naive australian_electricity 0.038 0.037 0.043 0.042 0.036 0.045 0.080 0.035 0.067 0.067 0.191 0.084 car_parts 1.010 1.013 1.005 1.000 0.995 1.004 1.054 0.975 1.652 1.333 1.340 1.600 cif_2016 0.017 0.017 0.016 0.017 0.016 0.018 0.014 0.014 0.053 0.033 0.039 0.015 covid_deaths 0.056 0.058 0.056 0.052 0.047 0.061 0.058 0.038 0.215 0.029 0.032 0.133 dominick 0.361 0.362 0.360 0.363 0.345 0.350 0.330 0.332 0.371 0.485 0.484 0.453 ercot 0.017 0.017 0.018 0.018 0.021 0.024 0.017 0.017 0.021 0.041 0.118 0.037 etth 0.075 0.075 0.079 0.079 0.071 0.077 0.076 0.090 0.085 0.089 0.132 0.122 ettm 0.062 0.059 0.072 0.067 0.052 0.051 0.071 0.073 0.052 0.105 0.085 0.141 exchange_rate 0.009 0.010 0.011 0.011 0.012 0.009 0.012 0.012 0.015 0.011 0.010 0.013 fred_md 0.036 0.047 0.048 0.042 0.042 0.033 0.027 0.052 0.027 0.035 0.033 0.122 hospital 0.054 0.054 0.054 0.054 0.057 0.058 0.055 0.053 0.050 0.059 0.053 0.073 m1_monthly 0.142 0.140 0.140 0.139 0.139 0.133 0.130 0.167 0.130 0.154 0.165 0.191 m1_quarterly 0.081 0.085 0.083 0.083 0.101 0.098 0.103 0.110 0.113 0.088 0.085 0.150 m1_yearly 0.133 0.133 0.161 0.161 0.151 0.162 0.185 0.126 0.145 0.133 0.139 0.209 m3_monthly 0.094 0.094 0.093 0.093 0.093 0.095 0.095 0.097 0.089 0.098 0.093 0.149 m3_quarterly 0.078 0.075 0.072 0.072 0.076 0.078 0.073 0.070 0.075 0.077 0.070 0.101 m3_yearly 0.126 0.124 0.148 0.142 0.129 0.157 0.147 0.131 0.144 0.156 0.129 0.167 m4_quarterly 0.079 0.077 0.076 0.075 0.077 0.078 0.082 0.076 0.062 0.079 0.079 0.119 m4_yearly 0.118 0.114 0.123 0.122 0.121 0.129 0.132 0.113 0.091 0.125 0.111 0.161 m5 0.563 0.561 0.562 0.561 0.562 0.566 0.584 0.572 0.557 0.617 0.628 1.024 nn5 0.190 0.190 0.188 0.190 0.150 0.157 0.157 0.147 0.155 0.248 0.264 0.425 nn5_weekly 0.086 0.084 0.085 0.085 0.084 0.085 0.090 0.090 0.079 0.084 0.088 0.123 tourism_monthly 0.100 0.100 0.096 0.095 0.090 0.098 0.095 0.104 0.085 0.091 0.100 0.104 tourism_quarterly 0.076 0.076 0.071 0.071 0.065 0.065 0.067 0.082 0.070 0.100 0.071 0.119 tourism_yearly 0.148 0.155 0.133 0.139 0.166 0.174 0.170 0.143 0.163 0.129 0.157 0.209 traffic 0.257 0.254 0.255 0.250 0.231 0.244 0.253 0.234 0.212 0.354 0.558 0.362 weather 0.134 0.134 0.135 0.132 0.134 0.130 0.138 0.132 0.133 0.215 0.215 0.217
Table 11: Mean absolute scaled errors (MASE) on Chronos Benchmark II
Chroma (4m, freq, best) Chroma (4m, freq, ens.) Chroma (tiny, freq, best) Chroma (tiny, freq, ens.) Chronos Bolt Base Chronos Bolt Mini Chronos Large Moirai Large TimesFM 2.0 Auto ARIMA Auto ETS Seasonal Naive australian_electricity 0.038 0.037 0.043 0.042 0.036 0.045 0.080 0.035 0.067 0.067 0.191 0.084 car_parts 1.010 1.013 1.005 1.000 0.995 1.004 1.054 0.975 1.652 1.333 1.340 1.600 cif_2016 0.017 0.017 0.016 0.017 0.016 0.018 0.014 0.014 0.053 0.033 0.039 0.015 covid_deaths 0.056 0.058 0.056 0.052 0.047 0.061 0.058 0.038 0.215 0.029 0.032 0.133 dominick 0.361 0.362 0.360 0.363 0.345 0.350 0.330 0.332 0.371 0.485 0.484 0.453 ercot 0.017 0.017 0.018 0.018 0.021 0.024 0.017 0.017 0.021 0.041 0.118 0.037 etth 0.075 0.075 0.079 0.079 0.071 0.077 0.076 0.090 0.085 0.089 0.132 0.122 ettm 0.062 0.059 0.072 0.067 0.052 0.051 0.071 0.073 0.052 0.105 0.085 0.141 exchange_rate 0.009 0.010 0.011 0.011 0.012 0.009 0.012 0.012 0.015 0.011 0.010 0.013 fred_md 0.036 0.047 0.048 0.042 0.042 0.033 0.027 0.052 0.027 0.035 0.033 0.122 hospital 0.054 0.054 0.054 0.054 0.057 0.058 0.055 0.053 0.050 0.059 0.053 0.073 m1_monthly 0.142 0.140 0.140 0.139 0.139 0.133 0.130 0.167 0.130 0.154 0.165 0.191 m1_quarterly 0.081 0.085 0.083 0.083 0.101 0.098 0.103 0.110 0.113 0.088 0.085 0.150 m1_yearly 0.133 0.133 0.161 0.161 0.151 0.162 0.185 0.126 0.145 0.133 0.139 0.209 m3_monthly 0.094 0.094 0.093 0.093 0.093 0.095 0.095 0.097 0.089 0.098 0.093 0.149 m3_quarterly 0.078 0.075 0.072 0.072 0.076 0.078 0.073 0.070 0.075 0.077 0.070 0.101 m3_yearly 0.126 0.124 0.148 0.142 0.129 0.157 0.147 0.131 0.144 0.156 0.129 0.167 m4_quarterly 0.079 0.077 0.076 0.075 0.077 0.078 0.082 0.076 0.062 0.079 0.079 0.119 m4_yearly 0.118 0.114 0.123 0.122 0.121 0.129 0.132 0.113 0.091 0.125 0.111 0.161 m5 0.563 0.561 0.562 0.561 0.562 0.566 0.584 0.572 0.557 0.617 0.628 1.024 nn5 0.190 0.190 0.188 0.190 0.150 0.157 0.157 0.147 0.155 0.248 0.264 0.425 nn5_weekly 0.086 0.084 0.085 0.085 0.084 0.085 0.090 0.090 0.079 0.084 0.088 0.123 tourism_monthly 0.100 0.100 0.096 0.095 0.090 0.098 0.095 0.104 0.085 0.091 0.100 0.104 tourism_quarterly 0.076 0.076 0.071 0.071 0.065 0.065 0.067 0.082 0.070 0.100 0.071 0.119 tourism_yearly 0.148 0.155 0.133 0.139 0.166 0.174 0.170 0.143 0.163 0.129 0.157 0.209 traffic 0.257 0.254 0.255 0.250 0.231 0.244 0.253 0.234 0.212 0.354 0.558 0.362 weather 0.134 0.134 0.135 0.132 0.134 0.130 0.138 0.132 0.133 0.215 0.215 0.217
Table 12: Weighted quantile losses (WQL) on GIFT-Eval (Part 1)
Chroma (4m, freq, best) Chroma (4m, freq, ens.) Chroma (tiny, freq, best) Chroma (tiny, freq, ens.) Chronos Bolt Base Chronos Bolt Small Chronos Large TTM-R2 Zeroshot TabPFN-TS TimesFM 2.0 (500m) Auto ARIMA Auto ETS Seasonal Naive LOOP-SEATTLE-5T-48 0.064 0.062 0.077 0.065 0.055 0.055 0.070 0.068 0.052 0.051 0.081 0.083 0.081 LOOP-SEATTLE-5T-480 0.165 0.158 0.146 0.138 0.116 0.119 0.176 0.121 0.087 0.110 0.123 0.238 0.123 LOOP-SEATTLE-5T-720 0.139 0.131 0.133 0.125 0.129 0.125 0.143 0.125 0.094 0.114 0.137 0.240 0.137 LOOP-SEATTLE-D-30 0.052 0.052 0.049 0.049 0.044 0.045 0.045 0.101 0.044 0.041 0.078 0.058 0.131 LOOP-SEATTLE-H-48 0.075 0.076 0.072 0.072 0.065 0.066 0.066 0.095 0.064 0.059 0.108 0.157 0.108 LOOP-SEATTLE-H-480 0.086 0.084 0.082 0.083 0.076 0.082 0.084 0.104 0.065 0.067 0.154 0.339 0.206 LOOP-SEATTLE-H-720 0.085 0.083 0.083 0.082 0.076 0.082 0.081 0.097 0.063 0.066 0.193 0.426 0.245 M-DENSE-D-30 0.082 0.078 0.087 0.087 0.069 0.072 0.075 0.151 0.057 0.060 0.135 0.123 0.294 M-DENSE-H-48 0.149 0.149 0.145 0.142 0.125 0.133 0.137 0.225 0.154 0.139 0.281 0.242 0.281 M-DENSE-H-480 0.147 0.146 0.141 0.140 0.157 0.134 0.134 0.213 0.159 0.127 0.255 0.458 0.479 M-DENSE-H-720 0.166 0.166 0.154 0.173 0.170 0.146 0.139 0.222 0.164 0.127 0.270 0.588 0.552 SZ-TAXI-15T-48 0.205 0.205 0.204 0.204 0.202 0.203 0.236 0.268 0.215 0.199 0.309 0.291 0.309 SZ-TAXI-15T-480 0.250 0.240 0.244 0.241 0.244 0.246 0.267 0.270 0.245 0.229 0.351 —– 0.454 SZ-TAXI-15T-720 0.248 0.242 0.255 0.249 0.248 0.245 0.265 0.260 0.248 0.227 0.398 —– 0.554 SZ-TAXI-H-48 0.138 0.137 0.138 0.137 0.136 0.137 0.148 0.183 0.144 0.135 0.170 1.620 0.229 bitbrains-fast-storage-5T-48 0.448 0.441 0.435 0.435 0.454 0.435 0.465 0.596 0.456 0.447 1.210 1.210 1.210 bitbrains-fast-storage-5T-480 0.801 0.792 0.825 0.803 0.755 0.867 0.798 0.906 0.735 0.881 1.270 1.270 1.270 bitbrains-fast-storage-5T-720 0.734 0.735 0.738 0.726 0.748 0.753 0.723 0.939 0.760 0.908 1.290 1.290 1.290 bitbrains-fast-storage-H-48 0.731 0.699 0.686 0.671 0.774 0.589 0.633 0.926 0.591 0.688 0.844 —– 1.080 bitbrains-rnd-5T-48 0.441 0.441 0.430 0.435 0.438 0.453 0.506 0.605 0.455 0.461 1.100 1.100 1.100 bitbrains-rnd-5T-480 0.751 0.729 0.743 0.733 0.605 0.792 0.636 0.835 0.710 0.727 1.260 1.260 1.260 bitbrains-rnd-5T-720 0.704 0.730 0.862 0.810 0.756 0.756 1.070 0.779 0.730 0.706 1.290 1.290 1.290 bitbrains-rnd-H-48 0.796 0.764 0.706 0.685 0.624 0.623 0.661 1.060 0.637 0.649 0.874 1.300 1.300 bizitobs-application-60 0.037 0.026 0.031 0.021 0.054 0.035 0.032 0.058 0.031 0.014 0.035 0.048 0.035 bizitobs-application-600 0.106 0.095 0.085 0.082 0.104 0.085 0.095 0.126 0.070 0.033 0.042 0.161 0.042 bizitobs-application-900 0.119 0.110 0.094 0.093 0.109 0.092 0.084 0.144 0.088 0.057 0.973 0.269 0.973 bizitobs-l2c-5T-48 0.079 0.079 0.078 0.077 0.074 0.073 0.084 0.107 0.099 0.084 0.262 0.075 0.262 bizitobs-l2c-5T-480 0.467 0.449 0.426 0.417 0.445 0.462 0.496 0.540 0.345 0.529 0.530 0.355 0.530 bizitobs-l2c-5T-720 0.785 0.737 0.673 0.649 0.738 0.790 0.741 0.785 0.511 0.748 0.674 0.597 0.674 bizitobs-l2c-H-48 0.221 0.215 0.209 0.210 0.189 0.204 0.491 0.549 0.290 0.345 0.547 0.494 0.536 bizitobs-l2c-H-480 0.288 0.290 0.277 0.278 0.254 0.285 0.802 0.782 0.380 0.640 0.813 1.050 1.420 bizitobs-l2c-H-720 0.314 0.310 0.308 0.300 0.278 0.295 0.757 0.751 0.440 0.728 0.787 1.110 1.820 bizitobs-service-60 0.032 0.023 0.025 0.019 0.051 0.032 0.027 0.053 0.031 0.015 0.040 0.046 0.040 bizitobs-service-600 0.091 0.081 0.074 0.075 0.096 0.082 0.072 0.117 0.067 0.038 0.049 0.224 0.049 bizitobs-service-900 0.115 0.106 0.097 0.095 0.113 0.096 0.094 0.140 0.091 0.062 0.056 0.373 0.056 car-parts-with-missing-12 1.010 1.013 1.005 1.000 0.995 1.007 1.070 2.285 0.955 1.046 1.290 1.340 1.720 covid-deaths-30 0.056 0.058 0.056 0.052 0.047 0.043 0.042 0.123 0.040 0.062 0.030 0.033 0.125 electricity-15T-48 0.093 0.092 0.091 0.090 0.082 0.082 0.092 0.152 0.104 0.079 0.165 0.175 0.165 electricity-15T-480 0.092 0.088 0.088 0.086 0.083 0.087 0.094 0.142 0.092 0.080 0.124 0.124 0.124 electricity-15T-720 0.092 0.089 0.092 0.087 0.084 0.086 0.094 0.143 0.089 0.083 0.129 0.129 0.129 electricity-D-30 0.058 0.058 0.056 0.056 0.055 0.058 0.061 0.093 0.060 0.060 0.083 0.111 0.122 electricity-H-48 0.075 0.074 0.073 0.073 0.064 0.067 0.065 0.097 0.072 0.054 0.109 0.099 0.109 electricity-H-480 0.095 0.094 0.091 0.088 0.081 0.084 0.086 0.109 0.091 0.073 0.156 0.156 0.156 electricity-H-720 0.110 0.108 0.102 0.109 0.098 0.102 0.107 0.128 0.112 0.089 0.190 —– 0.190 electricity-W-8 0.120 0.089 0.114 0.086 0.047 0.048 0.049 0.159 0.051 0.049 0.100 0.098 0.099 ett1-15T-48 0.173 0.170 0.171 0.169 0.158 0.169 0.196 0.235 0.183 0.168 0.241 0.383 0.241 ett1-15T-480 0.290 0.282 0.275 0.276 0.281 0.288 0.376 0.333 0.248 0.278 0.352 0.867 0.352 ett1-15T-720 0.289 0.284 0.306 0.295 0.298 0.296 0.367 0.352 0.260 0.283 0.396 1.050 0.396 ett1-D-30 0.311 0.311 0.295 0.273 0.287 0.283 0.404 0.416 0.292 0.281 0.279 0.334 0.515 ett1-H-48 0.197 0.195 0.193 0.192 0.181 0.189 0.198 0.250 0.194 0.192 0.223 0.440 0.250
Table 13: Weighted quantile losses (WQL) on GIFT-Eval (Part 2)
Chroma (4m, freq, best) Chroma (4m, freq, ens.) Chroma (tiny, freq, best) Chroma (tiny, freq, ens.) Chronos Bolt Base Chronos Bolt Small Chronos Large TTM-R2 Zeroshot TabPFN-TS TimesFM 2.0 (500m) Auto ARIMA Auto ETS Seasonal Naive ett1-H-480 0.309 0.318 0.298 0.343 0.303 0.295 0.333 0.339 0.276 0.282 0.384 —– 0.540 ett1-H-720 0.592 0.541 0.305 0.467 0.311 0.337 0.350 0.342 0.290 0.310 0.430 —– 0.616 ett1-W-8 0.281 0.253 0.277 0.260 0.296 0.293 0.316 0.448 0.256 0.272 0.305 0.277 0.338 ett2-15T-48 0.071 0.068 0.064 0.064 0.067 0.070 0.072 0.093 0.073 0.065 0.096 0.103 0.096 ett2-15T-480 0.119 0.115 0.109 0.107 0.110 0.119 0.126 0.128 0.098 0.105 0.143 0.325 0.143 ett2-15T-720 0.121 0.117 0.117 0.113 0.111 0.118 0.140 0.126 0.101 0.106 0.165 0.317 0.165 ett2-D-30 0.096 0.099 0.096 0.104 0.094 0.091 0.091 0.119 0.129 0.108 0.125 0.158 0.205 ett2-H-48 0.066 0.066 0.067 0.067 0.063 0.065 0.072 0.088 0.070 0.066 0.089 0.103 0.094 ett2-H-480 0.129 0.118 0.121 0.121 0.115 0.118 0.125 0.139 0.128 0.110 0.245 0.272 0.241 ett2-H-720 0.169 0.146 0.129 0.130 0.117 0.121 0.139 0.144 0.136 0.125 0.272 0.329 0.287 ett2-W-8 0.093 0.094 0.099 0.095 0.088 0.094 0.071 0.200 0.120 0.110 0.136 0.115 0.169 hierarchical-sales-D-30 0.577 0.576 0.580 0.577 0.576 0.582 0.599 0.792 0.593 0.576 0.735 0.931 2.360 hierarchical-sales-W-8 0.350 0.350 0.356 0.351 0.353 0.354 0.365 0.725 0.342 0.330 0.485 0.642 1.030 hospital-12 0.054 0.054 0.054 0.054 0.057 0.058 0.057 0.123 0.050 0.050 0.060 0.053 0.062 jena-weather-10T-48 0.035 0.034 0.046 0.048 0.033 0.037 0.044 0.045 0.035 0.016 0.155 0.082 0.155 jena-weather-10T-480 0.059 0.060 0.058 0.057 0.057 0.060 0.075 0.069 0.057 0.031 0.277 0.210 0.277 jena-weather-10T-720 0.060 0.059 0.070 0.067 0.064 0.063 0.077 0.068 0.055 0.035 0.304 0.282 0.304 jena-weather-D-30 0.046 0.046 0.046 0.046 0.045 0.047 0.049 0.124 0.047 0.058 0.080 0.074 0.297 jena-weather-H-48 0.043 0.046 0.042 0.044 0.042 0.043 0.046 0.060 0.043 0.045 0.143 0.158 0.173 jena-weather-H-480 0.055 0.055 0.059 0.059 0.054 0.058 0.071 0.073 0.060 0.066 0.211 —– 0.486 jena-weather-H-720 0.066 0.074 0.088 0.083 0.062 0.068 0.072 0.084 0.066 0.068 0.230 —– 0.598 kdd-cup-2018-with-missing-D-30 0.375 0.368 0.379 0.375 0.372 0.373 0.502 0.452 0.359 0.378 0.393 —– 0.888 kdd-cup-2018-with-missing-H-48 0.441 0.398 0.357 0.380 0.246 0.267 0.458 0.514 0.437 0.376 0.559 1.160 0.559 kdd-cup-2018-with-missing-H-480 0.456 0.443 0.423 0.425 0.301 0.364 0.658 0.532 0.462 0.466 0.851 0.949 0.949 kdd-cup-2018-with-missing-H-720 0.494 0.498 0.471 0.471 0.300 0.419 0.636 0.542 0.462 0.518 1.050 1.250 1.250 m4-daily-14 0.024 0.023 0.023 0.022 0.021 0.021 0.022 0.035 0.024 0.021 0.023 0.029 0.026 m4-hourly-48 0.028 0.029 0.027 0.026 0.025 0.020 0.026 0.040 0.028 0.011 0.034 0.070 0.040 m4-monthly-18 0.095 0.094 0.094 0.093 0.094 0.094 0.103 0.177 0.088 0.067 0.098 0.100 0.126 m4-quarterly-8 0.079 0.077 0.076 0.075 0.077 0.078 0.083 0.139 0.075 0.062 0.082 0.079 0.099 m4-weekly-13 0.043 0.044 0.043 0.043 0.038 0.038 0.037 0.069 0.036 0.042 0.050 0.052 0.073 m4-yearly-6 0.118 0.114 0.123 0.122 0.121 0.128 0.135 0.197 0.113 0.091 0.130 0.111 0.138 restaurant-30 0.287 0.282 0.279 0.278 0.264 0.264 0.279 0.438 0.297 0.261 0.362 —– 0.907 saugeenday-D-30 0.440 0.437 0.388 0.387 0.338 0.354 0.420 0.589 0.384 0.408 0.564 0.596 0.754 saugeenday-M-12 0.334 0.334 0.324 0.379 0.296 0.293 0.415 0.405 0.278 0.342 0.326 0.322 0.445 saugeenday-W-8 0.506 0.479 0.449 0.449 0.363 0.372 0.471 0.696 0.380 0.601 0.549 0.896 0.855 solar-10T-48 0.658 0.532 0.475 0.504 0.511 0.498 0.575 0.785 0.545 0.804 0.860 0.870 0.860 solar-10T-480 0.353 0.353 0.356 0.350 0.436 0.453 0.681 0.573 0.359 0.516 0.771 —– 0.771 solar-10T-720 0.355 0.355 0.340 0.334 0.443 0.497 0.747 0.545 0.352 0.498 0.786 2.930 0.786 solar-D-30 0.284 0.279 0.281 0.278 0.287 0.286 0.328 0.396 0.269 0.278 0.282 0.281 0.757 solar-H-48 0.322 0.322 0.317 0.316 0.298 0.303 0.337 0.468 0.358 0.406 0.628 1.080 0.628 solar-H-480 0.336 0.333 0.332 0.331 0.368 0.356 0.355 0.493 0.324 0.376 0.557 2.120 1.270 solar-H-720 0.341 0.339 0.338 0.343 0.405 0.373 0.448 0.512 0.324 0.493 0.607 1.720 1.470 solar-W-8 0.227 0.183 0.136 0.147 0.133 0.136 0.162 0.531 0.124 0.171 0.152 0.139 0.236 temperature-rain-with-missing-30 0.551 0.551 0.548 0.548 0.538 0.544 0.609 0.791 0.565 0.586 0.694 0.754 1.630 us-births-D-30 0.023 0.023 0.022 0.023 0.026 0.028 0.023 0.104 0.018 0.019 0.074 0.074 0.144 us-births-M-12 0.021 0.021 0.019 0.020 0.019 0.016 0.013 0.036 0.013 0.011 0.010 0.012 0.017 us-births-W-8 0.016 0.016 0.015 0.015 0.013 0.013 0.010 0.027 0.011 0.013 0.018 0.018 0.022
Table 14: Mean absolute scaled errors (MASE) on GIFT-Eval (Part 1)
Chroma (4m, freq, best) Chroma (4m, freq, ens.) Chroma (tiny, freq, best) Chroma (tiny, freq, ens.) Chronos Bolt Base Chronos Bolt Small Chronos Large TTM-R2 Zeroshot TabPFN-TS TimesFM 2.0 (500m) Auto ARIMA Auto ETS Seasonal Naive LOOP-SEATTLE-5T-48 0.064 0.062 0.077 0.065 0.055 0.055 0.070 0.068 0.052 0.051 0.081 0.083 0.081 LOOP-SEATTLE-5T-480 0.165 0.158 0.146 0.138 0.116 0.119 0.176 0.121 0.087 0.110 0.123 0.238 0.123 LOOP-SEATTLE-5T-720 0.139 0.131 0.133 0.125 0.129 0.125 0.143 0.125 0.094 0.114 0.137 0.240 0.137 LOOP-SEATTLE-D-30 0.052 0.052 0.049 0.049 0.044 0.045 0.045 0.101 0.044 0.041 0.078 0.058 0.131 LOOP-SEATTLE-H-48 0.075 0.076 0.072 0.072 0.065 0.066 0.066 0.095 0.064 0.059 0.108 0.157 0.108 LOOP-SEATTLE-H-480 0.086 0.084 0.082 0.083 0.076 0.082 0.084 0.104 0.065 0.067 0.154 0.339 0.206 LOOP-SEATTLE-H-720 0.085 0.083 0.083 0.082 0.076 0.082 0.081 0.097 0.063 0.066 0.193 0.426 0.245 M-DENSE-D-30 0.082 0.078 0.087 0.087 0.069 0.072 0.075 0.151 0.057 0.060 0.135 0.123 0.294 M-DENSE-H-48 0.149 0.149 0.145 0.142 0.125 0.133 0.137 0.225 0.154 0.139 0.281 0.242 0.281 M-DENSE-H-480 0.147 0.146 0.141 0.140 0.157 0.134 0.134 0.213 0.159 0.127 0.255 0.458 0.479 M-DENSE-H-720 0.166 0.166 0.154 0.173 0.170 0.146 0.139 0.222 0.164 0.127 0.270 0.588 0.552 SZ-TAXI-15T-48 0.205 0.205 0.204 0.204 0.202 0.203 0.236 0.268 0.215 0.199 0.309 0.291 0.309 SZ-TAXI-15T-480 0.250 0.240 0.244 0.241 0.244 0.246 0.267 0.270 0.245 0.229 0.351 —– 0.454 SZ-TAXI-15T-720 0.248 0.242 0.255 0.249 0.248 0.245 0.265 0.260 0.248 0.227 0.398 —– 0.554 SZ-TAXI-H-48 0.138 0.137 0.138 0.137 0.136 0.137 0.148 0.183 0.144 0.135 0.170 1.620 0.229 bitbrains-fast-storage-5T-48 0.448 0.441 0.435 0.435 0.454 0.435 0.465 0.596 0.456 0.447 1.210 1.210 1.210 bitbrains-fast-storage-5T-480 0.801 0.792 0.825 0.803 0.755 0.867 0.798 0.906 0.735 0.881 1.270 1.270 1.270 bitbrains-fast-storage-5T-720 0.734 0.735 0.738 0.726 0.748 0.753 0.723 0.939 0.760 0.908 1.290 1.290 1.290 bitbrains-fast-storage-H-48 0.731 0.699 0.686 0.671 0.774 0.589 0.633 0.926 0.591 0.688 0.844 —– 1.080 bitbrains-rnd-5T-48 0.441 0.441 0.430 0.435 0.438 0.453 0.506 0.605 0.455 0.461 1.100 1.100 1.100 bitbrains-rnd-5T-480 0.751 0.729 0.743 0.733 0.605 0.792 0.636 0.835 0.710 0.727 1.260 1.260 1.260 bitbrains-rnd-5T-720 0.704 0.730 0.862 0.810 0.756 0.756 1.070 0.779 0.730 0.706 1.290 1.290 1.290 bitbrains-rnd-H-48 0.796 0.764 0.706 0.685 0.624 0.623 0.661 1.060 0.637 0.649 0.874 1.300 1.300 bizitobs-application-60 0.037 0.026 0.031 0.021 0.054 0.035 0.032 0.058 0.031 0.014 0.035 0.048 0.035 bizitobs-application-600 0.106 0.095 0.085 0.082 0.104 0.085 0.095 0.126 0.070 0.033 0.042 0.161 0.042 bizitobs-application-900 0.119 0.110 0.094 0.093 0.109 0.092 0.084 0.144 0.088 0.057 0.973 0.269 0.973 bizitobs-l2c-5T-48 0.079 0.079 0.078 0.077 0.074 0.073 0.084 0.107 0.099 0.084 0.262 0.075 0.262 bizitobs-l2c-5T-480 0.467 0.449 0.426 0.417 0.445 0.462 0.496 0.540 0.345 0.529 0.530 0.355 0.530 bizitobs-l2c-5T-720 0.785 0.737 0.673 0.649 0.738 0.790 0.741 0.785 0.511 0.748 0.674 0.597 0.674 bizitobs-l2c-H-48 0.221 0.215 0.209 0.210 0.189 0.204 0.491 0.549 0.290 0.345 0.547 0.494 0.536 bizitobs-l2c-H-480 0.288 0.290 0.277 0.278 0.254 0.285 0.802 0.782 0.380 0.640 0.813 1.050 1.420 bizitobs-l2c-H-720 0.314 0.310 0.308 0.300 0.278 0.295 0.757 0.751 0.440 0.728 0.787 1.110 1.820 bizitobs-service-60 0.032 0.023 0.025 0.019 0.051 0.032 0.027 0.053 0.031 0.015 0.040 0.046 0.040 bizitobs-service-600 0.091 0.081 0.074 0.075 0.096 0.082 0.072 0.117 0.067 0.038 0.049 0.224 0.049 bizitobs-service-900 0.115 0.106 0.097 0.095 0.113 0.096 0.094 0.140 0.091 0.062 0.056 0.373 0.056 car-parts-with-missing-12 1.010 1.013 1.005 1.000 0.995 1.007 1.070 2.285 0.955 1.046 1.290 1.340 1.720 covid-deaths-30 0.056 0.058 0.056 0.052 0.047 0.043 0.042 0.123 0.040 0.062 0.030 0.033 0.125 electricity-15T-48 0.093 0.092 0.091 0.090 0.082 0.082 0.092 0.152 0.104 0.079 0.165 0.175 0.165 electricity-15T-480 0.092 0.088 0.088 0.086 0.083 0.087 0.094 0.142 0.092 0.080 0.124 0.124 0.124 electricity-15T-720 0.092 0.089 0.092 0.087 0.084 0.086 0.094 0.143 0.089 0.083 0.129 0.129 0.129 electricity-D-30 0.058 0.058 0.056 0.056 0.055 0.058 0.061 0.093 0.060 0.060 0.083 0.111 0.122 electricity-H-48 0.075 0.074 0.073 0.073 0.064 0.067 0.065 0.097 0.072 0.054 0.109 0.099 0.109 electricity-H-480 0.095 0.094 0.091 0.088 0.081 0.084 0.086 0.109 0.091 0.073 0.156 0.156 0.156 electricity-H-720 0.110 0.108 0.102 0.109 0.098 0.102 0.107 0.128 0.112 0.089 0.190 —– 0.190 electricity-W-8 0.120 0.089 0.114 0.086 0.047 0.048 0.049 0.159 0.051 0.049 0.100 0.098 0.099 ett1-15T-48 0.173 0.170 0.171 0.169 0.158 0.169 0.196 0.235 0.183 0.168 0.241 0.383 0.241 ett1-15T-480 0.290 0.282 0.275 0.276 0.281 0.288 0.376 0.333 0.248 0.278 0.352 0.867 0.352 ett1-15T-720 0.289 0.284 0.306 0.295 0.298 0.296 0.367 0.352 0.260 0.283 0.396 1.050 0.396 ett1-D-30 0.311 0.311 0.295 0.273 0.287 0.283 0.404 0.416 0.292 0.281 0.279 0.334 0.515 ett1-H-48 0.197 0.195 0.193 0.192 0.181 0.189 0.198 0.250 0.194 0.192 0.223 0.440 0.250
Table 15: Mean absolute scaled errors (MASE) on GIFT-Eval (Part 2)
Chroma (4m, freq, best) Chroma (4m, freq, ens.) Chroma (tiny, freq, best) Chroma (tiny, freq, ens.) Chronos Bolt Base Chronos Bolt Small Chronos Large TTM-R2 Zeroshot TabPFN-TS TimesFM 2.0 (500m) Auto ARIMA Auto ETS Seasonal Naive ett1-H-480 0.309 0.318 0.298 0.343 0.303 0.295 0.333 0.339 0.276 0.282 0.384 —– 0.540 ett1-H-720 0.592 0.541 0.305 0.467 0.311 0.337 0.350 0.342 0.290 0.310 0.430 —– 0.616 ett1-W-8 0.281 0.253 0.277 0.260 0.296 0.293 0.316 0.448 0.256 0.272 0.305 0.277 0.338 ett2-15T-48 0.071 0.068 0.064 0.064 0.067 0.070 0.072 0.093 0.073 0.065 0.096 0.103 0.096 ett2-15T-480 0.119 0.115 0.109 0.107 0.110 0.119 0.126 0.128 0.098 0.105 0.143 0.325 0.143 ett2-15T-720 0.121 0.117 0.117 0.113 0.111 0.118 0.140 0.126 0.101 0.106 0.165 0.317 0.165 ett2-D-30 0.096 0.099 0.096 0.104 0.094 0.091 0.091 0.119 0.129 0.108 0.125 0.158 0.205 ett2-H-48 0.066 0.066 0.067 0.067 0.063 0.065 0.072 0.088 0.070 0.066 0.089 0.103 0.094 ett2-H-480 0.129 0.118 0.121 0.121 0.115 0.118 0.125 0.139 0.128 0.110 0.245 0.272 0.241 ett2-H-720 0.169 0.146 0.129 0.130 0.117 0.121 0.139 0.144 0.136 0.125 0.272 0.329 0.287 ett2-W-8 0.093 0.094 0.099 0.095 0.088 0.094 0.071 0.200 0.120 0.110 0.136 0.115 0.169 hierarchical-sales-D-30 0.577 0.576 0.580 0.577 0.576 0.582 0.599 0.792 0.593 0.576 0.735 0.931 2.360 hierarchical-sales-W-8 0.350 0.350 0.356 0.351 0.353 0.354 0.365 0.725 0.342 0.330 0.485 0.642 1.030 hospital-12 0.054 0.054 0.054 0.054 0.057 0.058 0.057 0.123 0.050 0.050 0.060 0.053 0.062 jena-weather-10T-48 0.035 0.034 0.046 0.048 0.033 0.037 0.044 0.045 0.035 0.016 0.155 0.082 0.155 jena-weather-10T-480 0.059 0.060 0.058 0.057 0.057 0.060 0.075 0.069 0.057 0.031 0.277 0.210 0.277 jena-weather-10T-720 0.060 0.059 0.070 0.067 0.064 0.063 0.077 0.068 0.055 0.035 0.304 0.282 0.304 jena-weather-D-30 0.046 0.046 0.046 0.046 0.045 0.047 0.049 0.124 0.047 0.058 0.080 0.074 0.297 jena-weather-H-48 0.043 0.046 0.042 0.044 0.042 0.043 0.046 0.060 0.043 0.045 0.143 0.158 0.173 jena-weather-H-480 0.055 0.055 0.059 0.059 0.054 0.058 0.071 0.073 0.060 0.066 0.211 —– 0.486 jena-weather-H-720 0.066 0.074 0.088 0.083 0.062 0.068 0.072 0.084 0.066 0.068 0.230 —– 0.598 kdd-cup-2018-with-missing-D-30 0.375 0.368 0.379 0.375 0.372 0.373 0.502 0.452 0.359 0.378 0.393 —– 0.888 kdd-cup-2018-with-missing-H-48 0.441 0.398 0.357 0.380 0.246 0.267 0.458 0.514 0.437 0.376 0.559 1.160 0.559 kdd-cup-2018-with-missing-H-480 0.456 0.443 0.423 0.425 0.301 0.364 0.658 0.532 0.462 0.466 0.851 0.949 0.949 kdd-cup-2018-with-missing-H-720 0.494 0.498 0.471 0.471 0.300 0.419 0.636 0.542 0.462 0.518 1.050 1.250 1.250 m4-daily-14 0.024 0.023 0.023 0.022 0.021 0.021 0.022 0.035 0.024 0.021 0.023 0.029 0.026 m4-hourly-48 0.028 0.029 0.027 0.026 0.025 0.020 0.026 0.040 0.028 0.011 0.034 0.070 0.040 m4-monthly-18 0.095 0.094 0.094 0.093 0.094 0.094 0.103 0.177 0.088 0.067 0.098 0.100 0.126 m4-quarterly-8 0.079 0.077 0.076 0.075 0.077 0.078 0.083 0.139 0.075 0.062 0.082 0.079 0.099 m4-weekly-13 0.043 0.044 0.043 0.043 0.038 0.038 0.037 0.069 0.036 0.042 0.050 0.052 0.073 m4-yearly-6 0.118 0.114 0.123 0.122 0.121 0.128 0.135 0.197 0.113 0.091 0.130 0.111 0.138 restaurant-30 0.287 0.282 0.279 0.278 0.264 0.264 0.279 0.438 0.297 0.261 0.362 —– 0.907 saugeenday-D-30 0.440 0.437 0.388 0.387 0.338 0.354 0.420 0.589 0.384 0.408 0.564 0.596 0.754 saugeenday-M-12 0.334 0.334 0.324 0.379 0.296 0.293 0.415 0.405 0.278 0.342 0.326 0.322 0.445 saugeenday-W-8 0.506 0.479 0.449 0.449 0.363 0.372 0.471 0.696 0.380 0.601 0.549 0.896 0.855 solar-10T-48 0.658 0.532 0.475 0.504 0.511 0.498 0.575 0.785 0.545 0.804 0.860 0.870 0.860 solar-10T-480 0.353 0.353 0.356 0.350 0.436 0.453 0.681 0.573 0.359 0.516 0.771 4.590 0.771 solar-10T-720 0.355 0.355 0.340 0.334 0.443 0.497 0.747 0.545 0.352 0.498 0.786 2.930 0.786 solar-D-30 0.284 0.279 0.281 0.278 0.287 0.286 0.328 0.396 0.269 0.278 0.282 0.281 0.757 solar-H-48 0.322 0.322 0.317 0.316 0.298 0.303 0.337 0.468 0.358 0.406 0.628 1.080 0.628 solar-H-480 0.336 0.333 0.332 0.331 0.368 0.356 0.355 0.493 0.324 0.376 0.557 2.120 1.270 solar-H-720 0.341 0.339 0.338 0.343 0.405 0.373 0.448 0.512 0.324 0.493 0.607 1.720 1.470 solar-W-8 0.227 0.183 0.136 0.147 0.133 0.136 0.162 0.531 0.124 0.171 0.152 0.139 0.236 temperature-rain-with-missing-30 0.551 0.551 0.548 0.548 0.538 0.544 0.609 0.791 0.565 0.586 0.694 0.754 1.630 us-births-D-30 0.023 0.023 0.022 0.023 0.026 0.028 0.023 0.104 0.018 0.019 0.074 0.074 0.144 us-births-M-12 0.021 0.021 0.019 0.020 0.019 0.016 0.013 0.036 0.013 0.011 0.010 0.012 0.017 us-births-W-8 0.016 0.016 0.015 0.015 0.013 0.013 0.010 0.027 0.011 0.013 0.018 0.018 0.022
Appendix H Use of LLMs
Large language models are only used for a few word choices in the text. They are not used for the implementation of this research.